OpenStudy (anonymous):

Use the distributive property to rewrite the expression (a+b)4

OpenStudy (anonymous):

I'm assuming you mean \[(a+b)\times 4\] The law of distribution, which applies to multiplication, states that \[(a+b)\times c = (a \times c) + (b \times c)\]In your case c = 4.

OpenStudy (anonymous):

so if i well to rewrite it it would be (a+b)x4=(ax4)+(bx4)

OpenStudy (anonymous):

no i meant to write (a+8)4

OpenStudy (anonymous):

But do you mean: \[(a+8)\times 4\] or \[(a+8)^4\]Multiplication or exponentiation?

OpenStudy (anonymous):

(a+b)4

OpenStudy (anonymous):

If you mean multiplication then what you wrote is correct.